OpenStudy (anonymous):

What is the number of individual units (molecules or compounds) of Fe with a mass of 15.88 g (molar mass of 55 85g/mol) and has 0.28 moles?

OpenStudy (anonymous):

The number of particles in 0.28 moles is just 0.28* avogadro number= 0.28*6.022*10^23=1.686*10^23